OpenStudy (anonymous):

Find the vertex of the given parabola. Please show work. y=-x^2-2x+9

OpenStudy (anonymous):

vertex = -b/2a = -(-2) / 2(-2) = -1

OpenStudy (dumbcow):

^^ gives you the x-coordinate of vertex Substitute that value into equation to find y-coordinate

OpenStudy (anonymous):

The standard equation of a parabola is y = ax2 + bx + c. "vertex form": y = a(x – h)2 + k a= -1 b =2 -b/2a 2/2(-1) = -1 the x-coordinate of the vertex. Substituting in the original equation to get the y-coordinate, we get: y= -x^2 - 2x +9 y = (-1)^2 -2(-1) + 9 = 12 So, the vertex of the parabola is at (-1 , 12)

OpenStudy (dumbcow):

Watch the signs: -x^2 = -(x^2) x = -1 -(-1)^2 = -1 -1-2(-1) + 9 = -1 + 2 +9 = 10 vertex: (-1,10)